INSIGHT INTO THE JAHN-TELLER EFFECT IN THE NITRATE RADICAL VIA THE A-X HOT BANDS

\underline{KANA TAKEMATSU}, NATHAN EDDINGSAAS, and MITCHIO OKUMURA , Arthur Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena, California 91125 ; JOHN STANTON , Department of Chemistry, University of Texas at Austin, Austin, TX 78712.

Elucidation of the electronic states of the nitrate radical NO_{3} remains an interesting puzzle for experimentalists and theorists alike. The A^{2}E” state is Jahn-Teller active and is pseudo-Jahn-Teller coupled to the ground X^{2}A_{2}’ and excited B^{2}E’ states. Analysis of the vibronic-allowed A-X transition spectrum shows vibronic bands exhibiting both static and dynamic distortions depending on the vibrational level of the upper A state. In the E”xe’=a_{1}”+a_{2}”+e” manifold, only the a_{1}” levels are accessible under ambient temperature experiments. We have gained new insight into the A state by examining the hot bands of NO_{3}, as previously dark levels of the A state become accessible. We present analysis of our hot band spectrum along with calculations of the upper A state levels.